# Guided Transfer Learning for Discrete Diffusion Models

### v1

@@ D.1 Background
We can express the reverse process in terms of the forward transition rates and a concrete score defined by $s_{\theta}(\mathbf{y})_{\mathbf{y}^{\prime}}\approx\tfrac{q_{t}(\mathbf{y}^{\prime}\mid\mathbf{x})}{q_{t}(\mathbf{y}\mid\mathbf{x})}$, parameterized by a neural network $s_{\theta}:\mathcal{V}\to\mathbb{R}_{>0}^{|\mathcal{V}|}$. This enables training via concrete score matching (CSM) (meng2023concretescorematchinggeneralized), directly analogous to score-based models in continuous domains (song2021scorebasedgenerativemodelingstochastic). (lou2024discretediffusionmodelingestimating) propose the diffusion-weighted denoising score entropy (DWDSE) loss (their CDM-based objective) and note it can also be derived from the continuous-time likelihood framework of (campbell2022continuoustimeframeworkdiscrete); a detailed derivation appears in Appendix C.3 of (sahoo2024simpleeffectivemaskeddiffusion). Concretely,
@@
D.2 Theorem
In the previous section, we introduced the diffusion process and the source-domain objective on which the score function $s_{\theta}$ is trained. We now state a theorem that enables sampling from the target distribution without retraining $s_{\theta}$ under the objective in 37. Let $r_{\phi}(\mathbf{y})$ denote the learned density-ratio guidance term defined earlier.
@@ D.2 Theorem
D.3 Proof of Theorem 3
For the proof, we take the loss for a small step $\Delta t$ with $s=t-\Delta t$ from (36) and apply Theorem 1, which yields the following expression:
@@ D.3 Proof of Theorem 3
The work from (nisonoff2025unlockingguidancediscretestatespace) in equation (2) arrives at a similar term just for classifier-based guidance.

### v2

@@ F.1 Background
We can express the reverse process in terms of the forward transition rates and a concrete score defined by $s_{\theta}(\mathbf{y})_{\mathbf{y}^{\prime}}\approx\tfrac{q_{t}(\mathbf{y}^{\prime}\mid\mathbf{x})}{q_{t}(\mathbf{y}\mid\mathbf{x})}$, parameterized by a neural network $s_{\theta}:\mathcal{V}\to\mathbb{R}_{>0}^{|\mathcal{V}|}$. This enables training via concrete score matching (CSM) (Meng et al., 2023), directly analogous to score-based models in continuous domains (Song et al., 2021). Lou et al. (2024) propose the diffusion-weighted denoising score entropy (DWDSE) loss (their CDM-based objective) and note it can also be derived from the continuous-time likelihood framework of (Campbell et al., 2022); a detailed derivation appears in Appendix C.3 of (Sahoo et al., 2024). Concretely,
@@
F.2 Theorem
In the previous section, we introduced the diffusion process and the source-domain objective on which the score function $s_{\theta}$ is trained. We now state a theorem that enables sampling from the target distribution without retraining $s_{\theta}$ under the objective in 36. Let $r_{\phi}(\mathbf{y})$ denote the learned density-ratio guidance term defined earlier.
@@ F.2 Theorem
F.3 Proof of Theorem 3
For the proof, we take the loss for a small step $\Delta t$ with $s=t-\Delta t$ from equation 35 and apply Theorem 1, which yields the following expression:
@@ F.3 Proof of Theorem 3
The work from (Nisonoff et al., 2025) in equation (2) arrives at a similar term just for classifier-based guidance.